OpenStudy (anonymous):

3z + 7 ≤ 4z or 3z + 7 < -4z

OpenStudy (anonymous):

\[3z+7\leq 4z\] \[7\leq z\] OR \[3z+7< -4z\] \[7z<-7\] \[z<-1\]

OpenStudy (anonymous):

so either \[z<-1\] Or \[z\geq 7\] interval notation \[(-\infty<-1)\cup[7,\infty)\]